a sample of gas held in 2.7m3 at a pressure of 226kPa. The temperature is kept constant while the pressure is increased to 484kPa. Calculate the the new volume of the gas.


1
Expert's answer
2021-03-18T15:36:21-0400

"P_1 = 226kPa\\
V_1 = 2.7m³\\
P_2 = 484kPa\\
V_2 = \ ?"


from Boyle's law,

"{P_1V_1} = {P_2V_2}"

226 × 2.7 = 484 × V2


"V_2 = \dfrac{226 × 2.7 }{ 484}"


V2 = 1.26m³


The new volume of the gas is 1.26m³.